# Exploring nonlinearities in a positive ion-negative ion (PINI) plasma: can other processes mimic debris-induced effects?

### v3

@@ III.1.2 Comparison with pulse-induced nonlinearities
Figure 4: FCT simulation results of propagation of nonlinear wave in a PINI plasma with and embedded external negatively charged debris. These results are equivalent to the theoretical cases obtained through fKdV solution shown in Fig.2. The left panel (a) in the above figure corresponds to Fig.2(b) and the right panel (b) above corresponds to Fig.2(d). These results are plotted at the and of $\tau=100\equiv\varepsilon^{3/2}t$.
We argue that dynamically, an external perturbation can be considered to be having the same effect as an external charged debris so far as the leading edge (or precursor region) is concerned. So, we can expect a similar behavior of nonlinear wave front for this external perturbation and perturbation induced by an external debris. In this context, we would like to compare our results as shown in the cases for $\delta=1.02$ with $v_{d}=1.0,1.1$ in Fig.1(b) and (c) in the precursor regions and the case shown in Fig.4 of P25, which is comparable to our theoretical situation. In both cases, we have a leading shock front with oscillations which are nothing but DSWs when there is no dissipation. Note that Fig.4 of P25 is for the case when Landau damping is negligible with negligible negative ions $(\delta\sim 1.0)$. The situation is however different in the wake region (or trailing region of the pulse). This is due to the fact that experimentally a one-time external perturbation is induced by pushing a signal which then propagates in the form of discontinuity (either as a single soliton as a ramp structure) through the plasma which then transforms into a DSW as it moves. For the case of external debris however, the perturbation remains intact all throughout the life cycle of the nonlinear wave front giving rise to a very definitive signature in the wake region.
@@ III.2 Negatively charged external debris
We now present the results for a negatively charged debris in a flowing PINI plasma in Fig.2. All the parameters are kept same as in the case for positively charged debris. The shaded regions in all the panels show the extent of the charged debris with the centered dashed line indicating the debris position. As expected, bright pinned solitons are formed for $\delta$ close to unity at $\delta=1.02$, when the relative velocity between debris and PINI plasma $v_{d}$ exceeds a certain value as shown in Fig.2(a) and (b). It can be clearly seen that as $v_{d}$ increases, these pinned solitons finally vanishes even for $\delta\rightarrow 1$, which is consistent with our observations about pinned solitons in an $e$-$i$ plasma [17, 19]. For higher values of $\delta,$ IAWs in the precursor region propagate showing an oscillatory shock excitation which is not fully formed at a lower debris velocity. However as $v_{d}$ increases, amplitude of the precursor wave-front also increases and fully developed DSWs are observed to be formed. Here, it is worth mentioning that, similar behavior was also reported for the propagation of a rarefactive pulse in presence of positive ion beam in high negative ion density plasma [13]. Increasing the beam velocity resulted in a rise in soliton amplitude and a knee formation at the downstream, culminating in the formation of a shock-like structure. A realistic fluid simulation for much longer time period shows that these DSWs, as they propagate away from the site of the debris, get transformed into a train of individual solitons after segregation of the peaks, about which we shall discuss in Section IV. The pinned solitons at the site of the debris are deformed due to the presence of the shock. Also, behind the debris, soliton-like structures are seen to be excited due to ion-acoustic oscillations.
@@ IV Flux-corrected transport (FCT) simulation
Our FCT simulation results are presented through Fig.3 for the case with positively charged debris and through Fig.4 for negatively charged debris. In Fig.3, we have shown the results of the FCT simulation for the fKdV cases shown in Fig.1(c), corresponding to Fig.3(a) and Fig.1(d), corresponding to Fig.3(b). These curves are plotted at the end of $\tau=100\equiv\varepsilon^{3/2}t$. The $x$-axis is also respectively scaled for $\xi\equiv\varepsilon^{1/2}x$. Our estimate shows that the effective $\varepsilon\sim 0.7$ for $\rho_{0}=0.5$, which is used in the fKdV solution as well as in the FCT simulation. In both cases, the external charged debris is a Gaussian profile given by $\rho_{{\rm ext}}(\xi)=\rho_{0}e^{-\xi^{2}/\Delta}$. As we can see that the simulations results are quite comparable to those obtained from the corresponding fKdV solutions, except in the case for $\delta=1.5$ and $v_{d}=1.5$, which can be attributed to slightly different time over which various nonlinearities develop in a simulation and in fKdV solutions.
@@
It is important to note that the DSWs seen in Fig.2 and 4 are called ‘quasi-steady state shocks’ [39] as the initial oscillatory shock profile is not retained over a longer time period ($t\geq 200$). Fig.5 shows the potential distribution of one such shock profile at different time-steps $(t=0-100)$. Initially, the external charge perturbation introduces a highly nonlinear IAW and the leading edge of the precursor becomes steepened which forms a DSW with an oscillatory tail behind due to its dispersive nature. This non-steady shock profile, after a relatively long time $(t\geq 200)$, transforms into a train of individual propagating solitons as nonlinearity gets balanced by dispersion and each and every oscillatory peaks in the DSW separates into individual soliton structures. This effect is already observed and confirmed by earlier investigations on ion-acoustic shock waves both in theory and experiments [39]. The positive hump observed at the site of the debris is the signature formation of a single-peak bright pinned soliton.
@@ V Debris velocity and nonlinearity
One very interesting phenomenon is the changing property of the nonlinearity with increasing debris velocity which is similar to what is observed experimentally with respect to beam velocity in a PINI plasma [12]. From our analysis, what we observe is that as debris velocity increases, the wave amplitude steadily increases until reaching a maximum and then decreases, which is seen from the steepening in the leading edge (precursor region), which then gradually changes to a DSW before becoming a train of solitons and then a single KdV-like soliton. While the steepening in the leading edge indicates a dominant positive nonlinearity over dispersion at low debris velocity, as debris velocity increases, nonlinearity decreases and gets balanced by dispersion which creates a KdV-like soliton.
Figure 6: The changing nonlinearity with increasing debris velocity (from top to bottom) as obtained from the fKdV solutions (left column) and FCT simulation (right column). The thick, gray pulse in each frame is the debris-induced perturbation at the initial stage of the evolution which has later transformed into their respective form as time progresses.
In Fig.6, we show this effect as obtained from the fKdV solution [left column Fig.6(a-d)] and from FCT hydrodynamic simulation [right column Fig.6(e-h)]. In this figure, we have indicated the changing nonlinearity with an arrow with the keywords ‘steepening’, ‘DSW’, ‘train’, and ‘KdV’, respectively to indicate the phenomenon of steepening in the leading edge, formation of DSW, transformation to a train of solitons, and finally a simple KdV soliton as debris velocity $v_{d}$ increases from $0.8$ to $1.5$ for fKdV solutions and from $0.3$ to $1.8$ for FCT solutions. All the panels in the above figure are drawn at the rest frame of the debris, which is moving toward left. The thick, gray colored pulse in each frame indicates the initial stage of the perturbation induced by the moving debris which has later transformed. Density ratio $(\delta)$ and ion to electron temperature ratio $(\sigma)$ are taken to be $1.05$ and $0.05$ respectively, in both cases. It is important to note that all these evolutions are drawn at the same time evolution (column wise) across different $v_{d}$. The parameters are slightly different for fKdV and FCT solutions which are optimized respectively to represent this effect. This parameter difference of fKdV and FCT solutions can be attributed to the weak nonlinearity approximation that is inherent in the KdV formalism and full hydrodynamic simulation through FCT simulation.
@@ VI Summary and conclusion
To summarize, we have presented a detailed study of external debris-induced nonlinear structures in a multicomponent positive ion-negative ion (PINI) plasma. The motivation to choose a PINI plasma for this study is primarily due the fact that PINI plasmas are easily produced in low-temperature laboratory devices and also are widely used in plasma processing experiments. Additionally, theoretical and experimental (laboratory and also space-based) studies of debris-induced nonlinear plasma phenomena have been recently generating a lot of interests among the plasma physicists, especially due to the renewed attention of the scientific community in detection and removal of space-debris from low-earth orbits. Besides a detailed analysis of nonlinear waves excited by external charged debris, we also point out the similarities in certain characteristics with those excited by a high velocity positive ion beam and large amplitude pulse perturbations in such plasmas. The important findings of our analysis are presented below.
@@
  • For positively charged debris moving with subsonic velocities, small amplitude bright solitons are observed with steepening in the leading edge in low $\delta$ (low negative ion density) plasmas. On the other hand, steepening starts on the falling edge of the wave as $\delta$ is increased beyond certain value and as debris velocity reaches sonic and supersonic regime, large amplitude dark pinned solitons are excited. However, for the same velocity regime, DSWs are seen to be formed in low $\delta$ plasma. At very high debris velocity, the precursor wave is found to transform into a KdV like soliton.
@@
  • Toward the end, we have also shown how the KdV nonlinearity gets modified by the nonlinearity introduced by the external debris as debris velocity increases. This observation also has the similarity with the phenomenon of changing property of the nonlinear wave resulting out of beam-plasma interaction as beam velocity increases [11, 12, 13].
@@
We have argued that a high velocity ion beam would behave analogous to an moving external charge debris as long as the ion-acoustic timescale is smaller than the electron-ion collision time $(\tau_{{\rm IA}}\ll\tau_{ei})$ or beam ions move faster without remaining in phase with plasma waves, thus hindering beam plasma energy transfer process and would affect nonlinear wave dynamics via the electric field, similar to an external charge perturbation.
  • Toward this, we observe a resemblance between the DSWs excited by a moving debris in a low negative ion density PINI plasma to the oscillatory IA shocks excited by positive ramp signal in absence of the negative ions [24].
@@
  • Further the amplitudes of nonlinear waves are observed to increase for both positive and negative perturbation with increase in the debris velocity, which is similar to the case of increasing beam velocity in PINI plasma [11, 12, 13].
This work advances the understanding of externally driven nonlinear wave dynamics in multicomponent plasmas, offering a broader perspective on the applicability and limitations of perturbative approaches. The observed similarities between debris-driven nonlinear structures and and those due to other processes provide insights into the underlying mechanisms and may help in the interpretation of future experimental results involving negative ion plasmas.
@@ Acknowledgements.
Appendix A
@@
$\partial_{\tau}\phi^{(1)}+\frac{\beta}{\alpha}\phi^{(1)}\partial_{\xi}\phi^{(1)}+\frac{1}{\alpha}\partial_{\xi}^{3}\phi^{(1)}+\tilde{\gamma}\cdot{\cal P}\int_{-\infty}^{+\infty}\frac{\partial\phi^{(1)}}{\partial\xi^{\prime}}\frac{d\xi^{\prime}}{\xi-\xi^{\prime}}=0,$ (A1)
@@
$\lambda$ $=$ $1+\frac{3}{2}\sigma,$ (A2)
$\alpha$ $=$ $2\lambda^{-2}-3\lambda^{-4}\sigma,$ (A3)
$\tilde{\gamma}$ $=$ $\frac{\lambda}{\alpha\varepsilon\sqrt{2\pi}}\left[\left(\frac{m_{e}}{m_{+}}\right)^{1/2}+\sigma^{-3/2}\exp\left(-\frac{\lambda^{2}}{2\sigma}\right)\right],$ (A4)
and ${\cal P}$ denotes the principal value integration. The last term with the principal value integration in Eq.(A1) contains the linear Landau damping term of associated IAW and the Eq.(A1) as a whole accounts for weakly nonlinear IAW with Landau damping. We note that integration in Eq.(A1) is in the form of a convolution integral (with of course the principal value) and can be written as
${\cal P}\int_{-\infty}^{+\infty}\left(\frac{\partial\phi}{\partial\xi^{\prime}}\right)\frac{d\xi^{\prime}}{\xi-\xi^{\prime}}={\cal P}\left(\frac{1}{\xi}\right)*\left(\frac{\partial\phi}{\partial\xi^{\prime}}\right),$ (A5)
@@
${\cal F}\left[{\cal P}\int_{-\infty}^{+\infty}\left(\frac{\partial\phi}{\partial\xi^{\prime}}\right)\frac{d\xi^{\prime}}{\xi-\xi^{\prime}}\right]={\cal F}\left[{\cal P}\left(\frac{1}{\xi}\right)\right]{\cal F}\left[\left(\frac{\partial\phi}{\partial\xi^{\prime}}\right)\right].$ (A6)
@@
${\cal F}\left[{\cal P}\left(\frac{1}{\xi}\right)\right]$ $=$ $-i\pi\cdot{\rm sgn}(k),$ (A7)
${\cal F}\left[\left(\frac{\partial\phi}{\partial\xi^{\prime}}\right)\right]$ $=$ $ik\,\hat{\phi}(k),$ (A8)
@@
Considering now only the contribution of the Landau damping term, neglecting the nonlinear and the dispersion terms of Eq.(A1), and taking Fourier transform in space and time domain, we obtain the linear dispersion relation with only the Landau damping term as
$\omega=-i\cdot\tilde{\gamma}\cdot\pi k\cdot{\rm sgn}(k)\equiv-i\pi\tilde{\gamma}|k|,$ (A9)
@@
$\left[\tilde{\gamma}\cdot{\cal P}\int_{-\infty}^{+\infty}\frac{\partial\phi^{(1)}}{\partial\xi^{\prime}}\frac{d\xi^{\prime}}{\xi-\xi^{\prime}}\right]=\left[\nu\frac{\partial^{2}\phi^{(1)}}{\partial\xi^{2}}\right],$ (A10)
where the term on the right hand side is the viscous damping term in a KdV-Burgers equation. Note that Eq.(A1) is same as the KdV-Burgers equation with last term replaced by the dissipative damping term. The viscous damping rate of a fluid modeled by the KdV-Burgers equation is given by $(\nu k^{2})$, with $k$ being the wave number of the associated wave, which is in this case is IAW. So, by dimensional analysis, we have
$[\nu k^{2}]\simeq[\tilde{\gamma}k].$ (A11)

### v4

@@ III.1.2 Comparison with pulse-induced nonlinearities
Figure 4: FCT simulation results of propagation of nonlinear wave in a PINI plasma with an embedded external negatively charged debris. These results are equivalent to the theoretical cases obtained through fKdV solution shown in Fig.2. The left panel (a) in the above figure corresponds to Fig.2(b) and the right panel (b) above corresponds to Fig.2(d). These results are plotted at the end of $\tau=100\equiv\varepsilon^{3/2}t$.
We argue that dynamically, an external perturbation can be considered to be having the same effect as an external charged debris so far as the leading edge (or precursor region) is concerned. So, we can expect a similar behavior of nonlinear wave front for this external perturbation and perturbation induced by an external debris. In this context, we would like to compare our results as shown in the cases for $\delta=1.02$ with $v_{d}=1.0,1.1$ in Fig.1(b) and (c) in the precursor regions and the case shown in Fig.4 of P25, which is comparable to our theoretical situation. In both cases, we have a leading shock front with oscillations which are nothing but DSWs when there is no dissipation. Note that Fig.4 of P25 is for the case when Landau damping is negligible with negligible negative ions $(\delta\sim 1.0)$. The situation is however different in the wake region (or trailing region of the pulse). This is due to the fact that experimentally a one-time external perturbation is induced by pushing a signal which then propagates in the form of discontinuity (either as a single soliton-like pulse or a ramp structure) through the plasma which then transforms into a DSW as it moves. For the case of external debris however, the perturbation remains intact all throughout the life cycle of the nonlinear wave front giving rise to a very definitive signature in the wake region.
@@ III.2 Negatively charged external debris
We now present the results for a negatively charged debris in a flowing PINI plasma in Fig.2. All the parameters are kept same as in the case for positively charged debris. The shaded regions in all the panels show the extent of the charged debris with the centered dashed line indicating the debris position. As expected, bright pinned solitons are formed for $\delta$ close to unity at $\delta=1.02$, when the relative velocity between debris and PINI plasma $v_{d}$ exceeds a certain value as shown in Fig.2(a) and (b). It can be clearly seen that as $v_{d}$ increases, these pinned solitons finally vanish even for $\delta\rightarrow 1$, which is consistent with our observations about pinned solitons in an $e$-$i$ plasma [17, 19]. For higher values of $\delta,$ IAWs in the precursor region propagate showing an oscillatory shock excitation which is not fully formed at a lower debris velocity. However as $v_{d}$ increases, amplitude of the precursor wave-front also increases and fully developed DSWs are observed to be formed. Here, it is worth mentioning that, similar behavior was also reported for the propagation of a rarefactive pulse in presence of positive ion beam in high negative ion density plasma [13], where increasing the beam velocity resulted in a rise in soliton amplitude and a knee formation at the downstream, culminating in the formation of a shock-like structure. A realistic fluid simulation for much longer time period shows that these DSWs, as they propagate away from the site of the debris, get transformed into a train of solitary waves after segregation of the peaks, about which we shall discuss in Section IV. The pinned solitons at the site of the debris are deformed due to the presence of the shock. Also, behind the debris, soliton-like structures are seen to be excited due to ion-acoustic oscillations.
@@ IV Flux-corrected transport (FCT) simulation
Our FCT simulation results are presented through Fig.3 for the case with positively charged debris and through Fig.4 for negatively charged debris. In Fig.3, we have shown the results of the FCT simulation for the fKdV cases shown in Fig.1(c), corresponding to Fig.3(a) and Fig.1(d), corresponding to Fig.3(b). These curves are plotted at the end of $\tau=100\equiv\varepsilon^{3/2}t$. The $x$-axis is also respectively scaled for $\xi\equiv\varepsilon^{1/2}x$. Our estimate shows that the effective $\varepsilon\sim 0.7$ for $\rho_{0}=0.5$, which is used in the fKdV solution as well as in the FCT simulation. In both cases, the external charged debris is a Gaussian profile given by $\rho_{{\rm ext}}(\xi)=\rho_{0}e^{-\xi^{2}/\Delta}$. As we can see that the simulation results are quite comparable to those obtained from the corresponding fKdV solutions, except in the case for $\delta=1.5$ and $v_{d}=1.5$, which can be attributed to slightly different times over which various nonlinearities develop in a simulation and in fKdV solutions.
Figure 6: The changing nonlinearity with increasing debris velocity (from top to bottom) as obtained from the fKdV solutions (panels (a)-(d) correspond to $v_{d}=0.8,1.0,1.2$, and $1.5$, respectively) and FCT simulation (panels (e)-(h) correspond to $v_{d}=0.3,0.9,1.2$, and $1.8$, respectively). The thick, gray pulse in each frame is the debris-induced perturbation at the initial stage of the evolution which has later transformed into their respective form as time progresses.
@@
It is important to note that the DSWs seen in Fig.2 and 4 are called ‘quasi-steady state shocks’ [39] as the initial oscillatory shock profile is not retained over a longer time period ($t\geq 200$). Fig.5 shows the potential distribution of one such shock profile at different time-steps $(t=0-100)$. Initially, the external charge perturbation introduces a highly nonlinear IAW and the leading edge of the precursor becomes steepened which forms a DSW with an oscillatory tail behind due to its dispersive nature. This non-steady shock profile, after a relatively long time $(t\geq 200)$, transforms into a train of individual propagating solitary waves as nonlinearity gets balanced by dispersion and each and every oscillatory peaks in the DSW separates into individual solitary wave structures. This effect is already observed and confirmed by earlier investigations on ion-acoustic shock waves both in theory and experiments [39]. The positive hump observed at the site of the debris is the signature formation of a single-peak bright pinned soliton.
@@ V Debris velocity and nonlinearity
One very interesting phenomenon is the changing property of the nonlinearity with increasing debris velocity which is similar to what is observed experimentally with respect to beam velocity in a PINI plasma [12]. From our analysis, what we observe is that as debris velocity increases, the wave amplitude steadily increases until reaching a maximum and then decreases, which is seen from the steepening in the leading edge (precursor region), which then gradually changes to a DSW before becoming a train of solitary waves and then a single KdV soliton-like structure. While the steepening in the leading edge indicates a dominant positive nonlinearity over dispersion at low debris velocity, as debris velocity increases, nonlinearity decreases and gets balanced by dispersion which creates a KdV soliton-like structure. In Fig.6, we show this effect as obtained from the fKdV solution [left column Fig.6(a-d)] and from FCT hydrodynamic simulation [right column Fig.6(e-h)]. In this figure, we have indicated the changing nonlinearity with an arrow with the keywords ‘steepening’, ‘DSW’, ‘train’, and ‘KdV’, respectively to indicate the phenomenon of steepening in the leading edge, formation of DSW, transformation to a train of solitary waves, and finally a simple KdV soliton-like pulse as debris velocity $v_{d}$ increases from $0.8$ to $1.5$ for fKdV solutions and from $0.3$ to $1.8$ for FCT solutions. All the panels in the above figure are drawn at the rest frame of the debris, which is moving toward left. The thick, gray colored pulse in each frame indicates the initial stage of the perturbation induced by the moving debris which has later transformed. Density ratio $(\delta)$ and ion to electron temperature ratio $(\sigma)$ are taken to be $1.05$ and $0.05$ respectively, in both cases. It is important to note that all these evolutions are drawn at the same time evolution (column wise) across different $v_{d}$. The parameters are slightly different for fKdV and FCT solutions which are optimized respectively to represent this effect. This parameter difference of fKdV and FCT solutions can be attributed to the weak nonlinearity approximation that is inherent in the KdV formalism and full hydrodynamic simulation through FCT simulation.
@@ VI Summary and conclusion
To summarize, we have presented a detailed study of external debris-induced nonlinear structures in a multicomponent positive ion-negative ion (PINI) plasma. The motivation to choose a PINI plasma for this study is primarily due to the fact that PINI plasmas are easily produced in low-temperature laboratory devices and also are widely used in plasma processing experiments. Additionally, theoretical and experimental (laboratory and also space-based) studies of debris-induced nonlinear plasma phenomena have been recently generating a lot of interests among the plasma physicists, especially due to the renewed attention of the scientific community in detection and removal of space-debris from low-earth orbits. Besides a detailed analysis of nonlinear waves excited by external charged debris, we also point out the similarities in certain characteristics with those excited by a high velocity positive ion beam and large amplitude pulse perturbations in such plasmas.
The nonlinear wave structures observed in this study are a combined result of several physical mechanisms, governed by the response to the external debris perturbation by the plasma. Nonlinear precursor and wake waves are generated and propagated as an effect of the plasma ions in the vicinity of the perturbation going through ion-acoustic oscillations. The nature of debris charge also has a prominent effect on the plasma response[22, 19]. For instance, a positively charged debris leads to the formation of an ion hole at the location of perturbation and at higher debris velocities, the excited precursor waves transform into oscillatory DSWs. Whereas, negative debris introduces IICSI forming phase-space vortices at the debris site, where some ions get trapped which leads to the formation of pinned solitons. However for a PINI plasma, the character of the excited nonlinear structures, whether compressive (bright) or rarefactive (dark), is intrinsically determined by the concentration of the negative ions present. When there is a very small amount of negative ions present, the dynamics gets completely dominated by positive ions, and in the presence of a positive debris, usual ion-hole formation occurs that leads to the bright DSW formation in the precursor region. However, in the presence of higher amount of negative ions, IICSI-like phenomena occurs for negative negative ions, where, negative ions get attracted to the positive debris leading to the formation of dark pinned solitons. Opposite phenomena occurs when the debris is considered to be negatively charged and structures also get modified accordingly. The important findings of our analysis are presented below.
@@
  • For positively charged debris moving with subsonic velocities, small amplitude bright solitary waves are observed with steepening in the leading edge in low $\delta$ (low negative ion density) plasmas. On the other hand, steepening starts on the falling edge of the wave as $\delta$ is increased beyond a certain value and as debris velocity reaches the sonic and supersonic regime, large amplitude dark pinned solitons are excited. However, for the same velocity regime, DSWs are seen to be formed in low $\delta$ plasma. At very high debris velocity, the precursor wave is found to transform into a KdV soliton-like structure.
@@
  • Toward the end, we have also shown how the KdV nonlinearity gets modified by the nonlinearity introduced by the external debris as debris velocity increases. This observation also has a similarity with the phenomenon of changing property of the nonlinear wave resulting out of beam-plasma interaction as beam velocity increases [11, 12, 13].
@@
We have argued that a high velocity ion beam would behave analogous to a moving external charge debris as long as the ion-acoustic timescale is smaller than the electron-ion collision time $(\tau_{{\rm IA}}\ll\tau_{ei})$ or beam ions move faster without remaining in phase with plasma waves, thus hindering beam plasma energy transfer process and would affect nonlinear wave dynamics via the electric field, similar to an external charge perturbation.
  • Toward this, we observe a resemblance between the DSWs excited by moving debris in a low negative ion density PINI plasma to the oscillatory IA shocks excited by positive ramp signal in absence of the negative ions [24].
@@
  • Further, the amplitudes of nonlinear waves are observed to increase for both positive and negative perturbation with increase in the debris velocity, which is similar to the case of increasing beam velocity in PINI plasma [11, 12, 13].
It this context, we would like to draw a parallel between the nonlinear steepening that we have observed as the debris velocity increases and behavior of ‘density-divergence’ found to exist in the propagation of solitary wave in a cold collision-free plasma along the magnetic field [40]. This threshold-driven transition - debris velocity in our work (See Figs.1 and 5) and Mach number in the latter [40] is fundamental to these nonlinear waves despite their widely different plasma conditions. While our work is a purely electrostatic one and driven by external charged debris, the other one [40] investigate solitary wave in a cold plasma along a magnetic field.
This work advances the understanding of externally driven nonlinear wave dynamics in multicomponent plasmas, offering a broader perspective on the applicability and limitations of perturbative approaches. The observed similarities between debris-driven nonlinear structures and those due to other processes provide insights into the underlying mechanisms and may help in the interpretation of future experimental results involving negative ion plasmas.
@@ Acknowledgements.
Appendix A: Weak nonlinearity and fKdV formalism
Going by the expansions used in the derivation of the fKdV equation, Eqs.(8-10), the expansion parameter at any instant of time can be written as
$\varepsilon\sim\max\left(\left|\phi^{(1)}\right|,\left|n_{\pm}^{(1)}\right|,\left|v_{\pm}^{(1)}\right|\right).$ (A1)
However, in our case, this condition seems to be satisfied only in the initial phase and as the solution evolves in time, it definitely goes out of the weak nonlinearity regime and enters a fully nonlinear territory, where the reductive perturbation theory is not valid strictly. And yet, the fully nonlinear numerical solution of the system as obtained from the mFCT code closely resembles that of the fKdV solution.
This can be understood from the global property of the KdV-type solutions where nonlinearity and dispersion tend to balance each other resulting in a soliton. In this case, despite the external force term due to the charged debris, the global solution of the fKdV equation still maintains an equilibration between the leading-order nonlinearity (the quadratic $\phi$ term) and the leading-order dispersion (the third derivative term) preserving the overall nonlinear structure, resembling the simulation results. Toward this, we would like to refer to Infeld and Rowland [41], who have shown that KdV-like solitons persist and accurately describe the full nonlinear system even when the perturbation amplitude is not very small and KdV-like models can work way beyond their formal weak nonlinearity (small $\varepsilon$) limit. Similarly, Boyd [42] explains that KdV-like equations often remain remarkably accurate even when $\varepsilon\sim{\cal O}(1)$, especially for localized initial conditions (viz. an external debris) and smooth evolution (smaller time step).
Appendix B: External charged debris as first order perturbation
The reductive perturbation theory employed to derive the classical KdV equation essentially stems out of the linear theory. For a simplified $e$-$i$ plasma with no external debris, the linear dispersion relation can be written as
$\omega^{2}=k^{2}\left(\frac{1}{1+k^{2}}+\sigma\right),$ (B1)
where $\omega$ and $k$ are the frequency and wave number of perturbation with the linear perturbation defined as $\sim e^{-i\omega t+ikx}$. Note that in the above equation, the same normalization is used as in Section II. With an external debris having charged density $\rho_{\rm ext}$, the dispersion relation becomes
$k^{2}\left(1+\frac{1}{k^{2}\sigma-\omega^{2}}\right)+1=\frac{\rho_{\rm ext}}{\phi_{1}},$ (B2)
where $\phi_{1}$ denotes the first order perturbation in plasma potential. This reduces to Eq.(B1) for $\rho_{\rm ext}=0$. So, one can readily see that a physically relevant and correct linear dispersion relation exists only when $\rho_{\rm ext}\to 0$ at the first order or $\rho_{\rm ext}$ itself is expressed in terms of first order potential $\phi_{1}$. However, physically as $\rho_{\rm ext}$ is a purely external terms with no relation to plasma parameters, the latter possibility is ruled out leaving the only option of setting $\rho_{\rm ext}=0$ at the first order.
As the reductive perturbation theory is a nonlinear extension of the linear theory, this is also the precise reason why the reductive perturbation theory necessarily requires all first order external perturbations to vanish and any such perturbations must exist only at ${\cal O}(2)$ or higher.
Appendix C: KdV equation with linear Landau damping
@@
$\partial_{\tau}\phi^{(1)}+\frac{\beta}{\alpha}\phi^{(1)}\partial_{\xi}\phi^{(1)}+\frac{1}{\alpha}\partial_{\xi}^{3}\phi^{(1)}+\tilde{\gamma}\cdot{\cal P}\int_{-\infty}^{+\infty}\frac{\partial\phi^{(1)}}{\partial\xi^{\prime}}\frac{d\xi^{\prime}}{\xi-\xi^{\prime}}=0,$ (C1)
@@
$\lambda$ $=$ $1+\frac{3}{2}\sigma,$ (C2)
$\alpha$ $=$ $2\lambda^{-2}-3\lambda^{-4}\sigma,$ (C3)
$\tilde{\gamma}$ $=$ $\frac{\lambda}{\alpha\varepsilon\sqrt{2\pi}}\left[\left(\frac{m_{e}}{m_{+}}\right)^{1/2}+\sigma^{-3/2}\exp\left(-\frac{\lambda^{2}}{2\sigma}\right)\right],$ (C4)
and ${\cal P}$ denotes the principal value integration. The last term with the principal value integration in Eq.(C1) contains the linear Landau damping term of associated IAW and the Eq.(C1) as a whole accounts for weakly nonlinear IAW with Landau damping. We note that integration in Eq.(C1) is in the form of a convolution integral (with of course the principal value) and can be written as
${\cal P}\int_{-\infty}^{+\infty}\left(\frac{\partial\phi}{\partial\xi^{\prime}}\right)\frac{d\xi^{\prime}}{\xi-\xi^{\prime}}={\cal P}\left(\frac{1}{\xi}\right)*\left(\frac{\partial\phi}{\partial\xi^{\prime}}\right),$ (C5)
@@
${\cal F}\left[{\cal P}\int_{-\infty}^{+\infty}\left(\frac{\partial\phi}{\partial\xi^{\prime}}\right)\frac{d\xi^{\prime}}{\xi-\xi^{\prime}}\right]={\cal F}\left[{\cal P}\left(\frac{1}{\xi}\right)\right]{\cal F}\left[\left(\frac{\partial\phi}{\partial\xi^{\prime}}\right)\right].$ (C6)
@@
${\cal F}\left[{\cal P}\left(\frac{1}{\xi}\right)\right]$ $=$ $-i\pi\cdot{\rm sgn}(k),$ (C7)
${\cal F}\left[\left(\frac{\partial\phi}{\partial\xi^{\prime}}\right)\right]$ $=$ $ik\,\hat{\phi}(k),$ (C8)
@@
Considering now only the contribution of the Landau damping term, neglecting the nonlinear and the dispersion terms of Eq.(C1), and taking Fourier transform in space and time domain, we obtain the linear dispersion relation with only the Landau damping term as
$\omega=-i\cdot\tilde{\gamma}\cdot\pi k\cdot{\rm sgn}(k)\equiv-i\pi\tilde{\gamma}|k|,$ (C9)
@@
$\left[\tilde{\gamma}\cdot{\cal P}\int_{-\infty}^{+\infty}\frac{\partial\phi^{(1)}}{\partial\xi^{\prime}}\frac{d\xi^{\prime}}{\xi-\xi^{\prime}}\right]=\left[\nu\frac{\partial^{2}\phi^{(1)}}{\partial\xi^{2}}\right],$ (C10)
where the term on the right hand side is the viscous damping term in a KdV-Burgers equation. Note that Eq.(C1) is same as the KdV-Burgers equation with last term replaced by the dissipative damping term. The viscous damping rate of a fluid modeled by the KdV-Burgers equation is given by $(\nu k^{2})$, with $k$ being the wave number of the associated wave, which is in this case is IAW. So, by dimensional analysis, we have
$[\nu k^{2}]\simeq[\tilde{\gamma}k].$ (C11)